If A atom is at corner B atom is at face centre and A atoms is missing from 2 corners in each unit cell what is simplest formula?

Grade:12

1 Answers

Vikas TU
7 years ago
Proportion of the each atom at the corner = > 1/8
For (8-2) corners => total number of atoms  => 6/8 = 3/4
For B  = 6*1/2 = 3
The formulae becomes => 3(A1/4B)